# labelinglabel

### v1

@@ 6.2 Kendall-$\tau$ rank correlation
Table 2: Kendall-$\tau$ rank correlation coefficient results
6.2.1 $\tau$-Correlation Result Analysis and Conclusions
@@
Our findings suggest that a $\tau$-correlation score of approximately $0.70$ is sufficient to distinguish between high- and low-quality LaaJs, even when model distance is only roughly estimated. This insight supports the use of $\tau$-correlation as a reliable metric and helps establish a practical threshold for LaaJ evaluation. We conclude that $\tau$-correlation is an effective and robust metric for LaaJ meta-evaluation.
6.3 Ordering Experiment
In our final example, we demonstrate sensitivity to distance and thereby learn to apply a given metric appropriately. A test we use for meta-evaluation is the ability of a LaaJ to correctly identify the better model among two candidates. To this end, we select two models of known, comparable quality, apply them to the same benchmarks, and compute the percentage of points for which the LaaJ (weakly) prefers the better model.
It is important to note that the better model does not necessarily produce superior inferences for all points. However, we expect that a more accurate LaaJ will yield higher results in this test, as it is less noisy, makes fewer mistakes, and is thus more likely to be biased toward the better model. We conducted a simulation of the ordering experiment using virtual models at various distances. The results are presented in Table 3.
@@ 6.3 Ordering Experiment
Table 3: Ordering experiment results
6.3.1 Ordering Experiment Result Analysis and Conclusions
@@
For example, a result of approximately $80\%$ is achieved by $L_{3}$ at distance 4, and also by the noisier $L_{10}$ at distance 10. This illustrates that without accounting for model distance, the metric may fail to distinguish between LaaJs of different quality.
@@
7 Conclusion
@@
Through a detailed use case in code translation, we demonstrated how LaaJMeter can simulate nuanced model and LaaJ behaviors. We evaluated several metrics, $t$-test, Kendall-$\tau$ correlation, and an ordering experiment, highlighting their strengths and limitations. Our findings show that while the $t$-test lacks sensitivity to LaaJ quality, Kendall-$\tau$ correlation provides a robust signal even with imprecise model distance estimates. The ordering experiment, although effective, requires careful calibration based on model distance.
In this paper, we focused on simulating evaluation metrics that do not rely on human-annotated data. However, our framework can be extended to incorporate such data. This extension would involve simulating expert scores. One straightforward approach is to define the ground-truth scores as human evaluations. Alternatively, noise can be added to the ground-truth scores to simulate variability in expert judgment. This enables the evaluation of human-referenced metrics, such as mean squared error. We leave the exploration of this direction for future work.

### v2

@@ VI-A $t$-Test
Table II: Kendall-$\tau$ rank correlation coefficient results
Importantly, the distance between the models also plays a critical role. If the distance is too large, even a poor LaaJ may detect the difference, leading to a significant $p$-value. To explore this, we simulated $t$-tests on virtual LaaJ scores across pairs of virtual models with varying distances. The resulting $p$-values are shown in Table I.
VI-A1 $t$-Test Result Analysis and Conclusions
As shown in Table I, all LaaJs yield statistically significant $p$-values for model distances greater than 2. This indicates that even a poor LaaJ, such as $L_{10}$, can detect differences between models, even when the distance is merely notable. Consequently, the $t$-test lacks sensitivity to LaaJ quality and fails to distinguish between LaaJs of varying performance. We conclude that the $t$-test is not suitable for meta-evaluation in this context.
VI-B Kendall-$\tau$ rank correlation
Moreover, we expect the difference in $\tau$-correlation between good and poor LaaJs to be more pronounced when comparing models with small distances. In such cases, subtle differences in inference quality are harder to detect, and only better LaaJs are expected to capture them effectively. The results of our simulation are presented in Table II.
VI-B1 $\tau$-Correlation Result Analysis and Conclusions
@@
Our findings suggest that a $\tau$-correlation score of approximately $0.70$ is sufficient to distinguish between high and low quality LaaJs, even when model distance is only roughly estimated. This insight supports the use of $\tau$-correlation as a reliable metric and helps establish a practical threshold for LaaJ evaluation. We conclude that $\tau$-correlation is an effective and robust metric for meta-evaluation of a C2J LaaJ.
VI-C Ordering Experiment
@@
Table III: Ordering experiment results
In our final example, we demonstrate sensitivity to distance and thereby learn to apply a given metric appropriately. A test we frequently use for meta-evaluation is the ability of a LaaJ to correctly identify the better model among two candidates. To this end, we select two models of known, comparable quality, apply them to the same benchmarks, and compute the percentage of points for which the LaaJ (weakly) prefers the better model.
It is important to note that the better model does not necessarily produce superior inferences for all points. However, we expect that a more accurate LaaJ will yield higher results in this test, as it is less noisy, makes fewer mistakes, and is thus more likely to be biased toward the better model. We conducted a simulation of the ordering experiment using virtual models at various distances. The results are presented in Table III.
VI-C1 Ordering Experiment Result Analysis and Conclusions
@@
For example, a result of approximately $80\%$ is achieved by $L_{3}$ at distance 4, and also by the noisier $L_{10}$ at distance 10. This illustrates that without accounting for model distance, the metric may fail to distinguish between LaaJs of sufficent quality.
@@
VII Distance Estimation
The results presented in Section VI imply that, for some metrics, estimating model distance is necessary. The ordering experiment for C2J LaaJs is one example of this phenomenon. To support such estimation, we added an automated distance estimation component to the ordering experiments we perform. By running the experiment across several candidate LaaJs and two models of known, differing quality, we compute three estimates of model distance. Furthermore, for a given LaaJ and distance estimate, the automation identifies the virtual LaaJ among $L_{1}\dots L_{10}$ whose performance is closest to that of the evaluated LaaJ (see Section V-B for the explicit construction).
The first distance estimate is the self-reference distance: for a C2J LaaJ, this is defined as the difference between the mean scores assigned by the LaaJ to the two models. The second estimate is the average distance: the average of all self-reference distances across the LaaJs used in the experiment. The third is the best performer distance: the self-reference distance of the LaaJ that achieved the highest score in the ordering experiment. That is, the one that assigned a higher score to the second model for the greatest number of points.
The three options for distance estimations are intuitive methods our team is currently exploring. We do not claim that these are optimal, nor have we developed techniques to determine which among them is preferable. Developing methods to validate distance estimations and implementing them is beyond the scope of this paper and is left for future research.
VIII Conclusion
@@
We reported a use case of using LaaJMeter for meta-evaluation of C2J LaaJs. A C2J laaj evaluates code translations of COBOL to Java, a service provided by IBM WCA4Z. Thtough the use case we demonstrated how LaaJMeter can simulate nuanced model and LaaJ behaviors. We evaluated several metrics, $t$-test, Kendall-$\tau$ correlation, and an ordering experiment, highlighting their strengths and limitations. Our findings show that while the $t$-test lacks sensitivity to LaaJ quality, Kendall-$\tau$ correlation provides a robust signal even with imprecise model distance estimates. The ordering experiment, although effective, requires careful calibration based on model distance.
VIII-A Future Research
LaaJMeter is a framework designed to support meta-evaluation in scenarios where annotated data is scarce. It achieves this by applying metrics to virtual models and LaaJs, enabling analysis of metrics’ ability to distinguish between high and low quality LaaJs. Essentially, the framework is valuable because it allows metric simulations to be run on synthetic data. This approach, using metric simulation to assess metric quality, is not exclusive to meta-evaluation. An important direction for future research is to generalize the construction of the virtual components (models and LaaJs, in our case) so that the core concepts of LaaJMeter can be applied to other domains as well.
A key factor in successful simulations using the LaaJMeter framework is the adequate construction of virtual models and LaaJs. For the C2J use case, we relied on experience, understanding, and intuition gained during the project to construct models that reflect reality as faithfully as possible. We wonder whether statistical tests could be developed to assess how well the virtual components reflect reality, for example, by observing LaaJ scores across various real benchmarks. Furthermore, it remains an open question whether virtual components could be constructed automatically, given LaaJ executions over benchmark datasets.
We noted that metrics may be sensitive to the distance between models, which we defined as the mean (ground truth) score difference for model outputs on the same input. Possible future research directions include validating whether a distance estimate is approximately correct and developing methods to automatically extract valid distance estimates. Since ground truth scores for models are not available, this task is inherently challenging. Moreover, precise distance estimation may be infeasible, and only approximate estimates can be expected.